OpenStudy (anonymous):

could someone check these answers for me please as the quiz i have just taken say my answers are incorrect: (1) A PERSON WANTS TO PLACE PAVERS FOR HIS PATIO: 4 2/3 YARDS by 5 1/3 Yards = 10? Yards? 9sq feet per Yard so 90 sq ft @ $3.35 =$301.50? (2) Average customer spends $67 at checkout, store has 4 registers open each handles 24 customers per hr estimate the amount of money the store receives in 1 hour this is the correct formula ? 67*4*24 = 6,432? Thanks in advance

OpenStudy (anonymous):

I'm confused on what you mean in 1, I'm still working on that. But for 2: 4 registers can handle 24 customers each, multiply that to find how many customers total come through per hour. Then each spends 67 dollars so you would multiply that there. You did the right formula. As long as the math is right it should be correct.

OpenStudy (anonymous):

For the first one I'm not sure what you did. This is what I get: First you find the area of the patio. (A=L*W) Area = 4 2/3 * 5 1/3 = 24.8378.... square yards. 9 sq ft per sq yard, so multiply by 9. = 223.54 sq ft.

OpenStudy (anonymous):

thanks for response, in regards to question 1 they quiz answer they say is $5600

OpenStudy (anonymous):

bang on with answer for patio, I added instead of multiplying my brain froze I guess math is correct on Q2 have sent in email for clarification